A system of equations is shown below.3x + 2y = 19 2x - y = 1 What is the solution to the system of equations?
Answer by checkley79(3341)

3x + 2y = 19
2x - y = 1 MULTIPLY BY 2 & ADD.
3X+2Y=19
4X-2Y=2
-----------------
7X=21
X=21/7
X=3 ANS.
3*3+2Y=19
9+2Y=19
2Y=19-9
2Y=10
Y=10/2
Y=5 ANS.
PROOF:
2*3-5=1
6-5=1
1=1